29. Consider the following thermochemical equations. Cu20(s) → Cu( s) + CuO(s) Δ1P +11 kJ Calculate the standard enthalpy of formation of CuO(s). a) 166 k.J d) +155 k.J b)-299 kJ e)--155 kJ c) +299 kJ

Please Answer and explain. thanks
Answer #1

Given that

Cu2O (s) + 1/2 O2 gives 2CuO   \DeltaH = -144 kJ equation 1

Cu2O (s) gives Cu + CuO \DeltaH = 11 kJ equation 2

reverse the direction (note: while reversing the direction of a reaction the sign of \DeltaH will change)

Cu + CuO gives Cu2O (s) \DeltaH = -11 kJ equation 3

add equation 3 with equation 1

Cu2O (s) + 1/2 O2 + Cu + CuO gives Cu2O (s) + 2CuO   \DeltaH = -155 kJ

cancel the similar terms on both side of the eqaution

then

1/2 O2 + Cu gives CuO \DeltaH = -155 kJ

option e is the correct answer.